at an instant car A Is moving with constant velocity of 30 m/s and a stationary car B with acc. 2 m/s^2 in same direction and location, the time after which they cross each other

Grade:9

1 Answers

Vikas TU
7 years ago
Let car A covered the distance as 'x’ meters then,
x = 30*t.................(1)
 
For car B,
x = 0 + 0.5*2*t^2...............(2)
solving both the eqns. we get,
t^2 = 30t
t (t – 30 ) = 0
t = 30 seconds.